a manufactrer produces lightbulbs that are packed into boxes of 100. if quality control studies indicate that 0.5% of lightbulbs produce are defective, what percentage of the boxes will contain if 2 or more defectives?


1
Expert's answer
2022-03-29T05:12:21-0400

As n is large and p, the P (defective bulb), is small, use the Poisson approximation to the Binomial probability distribution.

0.5%=0.005

"\lambda=100(0.005)=0.5"

"P(x \ge2)=1-P(x \le1)=1-0.9098=0.0902"